# Screened Thin-Target Bremsstrahlung with Partially-Ionized High-Z Species

### v1

@@ I Introduction
A recurring limitation of analytic bremsstrahlung treatments is the incorporation of screening. Classical prescriptions (Molière parameterizations [34], Thomas–Fermi–like screening lengths [52, 9], and tabulated atomic form factors [19, 49]) describe neutral-atom behavior but do not by themselves provide a compact analytic route to interpolate across charge states. Alternative analytic screening laws have also been proposed, such as the Thomas–Fermi–Kirillov form (TFK), which replaces the Yukawa exponential by a power $n=3/2$ [26], see Eq. (32), but these remain difficult to integrate consistently into bremsstrahlung cross-section formalisms. Model-potential approaches, most notably that of Avdonina and Pratt [1], provide analytic screened atomic form factors (AFF) and explicitly address ionic charge states, demonstrating that screening significantly alters spectral and polarization observables. Their framework, however, relies on a specific one-parameter potential and is not easily generalized to reproduce high-quality tabulations or to interpolate smoothly between neutral and bare-ion limits. At the same time, investigations of Coulomb corrections emphasize that screening and higher-order distortion must be treated together to achieve percent-level accuracy in the application-relevant $\sim$1–100 MeV window and for moderate to high-$Z$ targets [43, 32].
To overcome these difficulties, we introduce an analytic AFF representation built from a finite sum of Yukawa screening potentials [17, 50]. Each term admits a simple analytic momentum-space component suitable for insertion into differential cross sections, while the linear combination can be fitted to high-quality atomic data (Dirac–Hartree–Fock–Slater (DHFS) fits [49, 12, 11] or Hubbell tabulations [19]) and adjusted to account for removed bound charge, thereby extending from neutral atoms to fully stripped ions. Practically, screening is implemented by adding the fitted multi-Yukawa (MY) form factor into the Bethe–Heitler (BH) differential cross section while higher-order Coulomb distortion is retained through the Roche–Ducos–Proriol (RDP) expression for the unscreened contribution [43]. Following the Olsen–Maximon–Wergeland (OMW) additivity rule [37] (see Sec. II), the screened doubly differential cross section is obtained as the sum of the unscreened, exact expression in a pure Coulomb field, and a screening correction in the first Born approximation, such that both neutral-atom and bare-nucleus limits are preserved. The result is a compact, rapidly evaluable expression for the doubly differential cross section (DDCS) that is inexpensive compared with full PW models.
@@ II Theoretical Framework for Electron–Ion Bremsstrahlung
[…]
The triply-differential cross section (TDCS) in Eq. (1) involves an integral over all space of the initial and final electron wave functions. In the vicinity of the nucleus, the nuclear (Coulomb) effects are important, while further from it, the electrons bound to the nucleus may screen the potential of the latter, so that a screening correction to the exact Coulomb field cross section ought to be considered. Olsen, Maximon and Wergeland noted that for the DDCS, i.e. after integrating out the electron’s degrees of freedom, and for energies greater than the electron rest mass, the screening correction to the Coulomb cross section is additive [37]:
$\left(\frac{d^{2}\sigma_{e-i}^{Z_{s^{\prime},s}}}{dkd\Omega_{k}}\right)^{\text{s}}_{\text{exact}}\simeq\left(\frac{d^{2}\sigma_{e-i}^{Z_{s}}}{dkd\Omega_{k}}\right)_{\text{exact}}^{\text{us}}+\left[\left(\frac{d^{2}\sigma_{e-i}^{Z_{s^{\prime},s}}}{dkd\Omega_{k}}\right)_{\text{Born}}^{\text{s}}-\left(\frac{d^{2}\sigma_{e-i}^{Z_{s}}}{dkd\Omega_{k}}\right)_{\text{Born}}^{\text{us}}\right].$ (3)
The latter is the so-called OMW additivity rule. The term in brackets is the screening correction. The superscripts s and us mean screened and unscreened, respectively. An illustrating figure is presented in Appendix A. We now introduce the expressions we will consider for the terms from Eq. (3).
@@ II.1 Bethe–Heitler Cross Section
$\left(\frac{d^{3}\sigma_{e-i}^{Z_{s^{\prime},s}}}{d\bar{k}d\Omega_{k}d\Omega_{p}}\right)_{\text{BH}}=$ $\alpha\left(\frac{r_{e}}{2\pi}\right)^{2}Z_{s}^{2}\left(1-F_{Z_{s^{\prime},s}}\left(\bar{q}\right)\right)^{2}$ (4)
$\times\frac{\bar{p}}{\bar{k}\bar{p}_{0}\bar{q}^{4}}A^{\text{BH}},$ (5)
with $r_{e}$ the classical electron radius, and where $F_{Z_{s^{\prime},s}}$ is the form factor. The latter is set to zero in the unscreened case, and we retain this expression for the unscreened bremsstrahlung term, i.e. the third on the right-hand side of Eq. (3). The recoil momentum transferred to the nucleus, $\bar{q}$, is given by the relation
$\bar{q}^{2}=$ $\bar{p}_{0}^{2}+\bar{p}^{2}+\bar{k}^{2}-2\bar{p}_{0}\bar{k}\cos\theta_{0}+2\bar{p}\bar{k}\cos\theta$ (6)
$-2\bar{p}\bar{p}_{0}\left(\cos\theta\cos\theta_{0}+\sin\theta\sin\theta_{0}\cos\phi\right),$ (7)
with $\theta_{0}$ the angle between the incoming electron and the emitted photon. The barred quantities are normalized with respect to the electron rest-mass energy $m_{e}c^{2}$ or momentum $m_{e}c$. Here, and from now on, $Z_{s^{\prime},s}$ represents the effective ionic charge of species $Z_{s}$, that is, the atomic number minus the number of bound electrons. The term $A^{\text{BH}}$ in Eq. (5) is given
$A^{\text{BH}}=\sum_{n=1}^{n=4}A_{n}^{\text{BH}},$ (8)
@@
$A_{1}^{\text{BH}}$ $=\frac{\bar{p}^{2}\sin^{2}\theta}{\left(\bar{E}-\bar{p}\cos\theta\right)^{2}}\left(4\bar{E}_{0}^{2}-\bar{q}^{2}\right)$ (9a)
$A_{2}^{\text{BH}}$ $=\frac{\bar{p}_{0}^{2}\sin^{2}\theta_{0}}{\left(\bar{E}_{0}-\bar{p}_{0}\cos\theta_{0}\right)^{2}}\left(4\bar{E}^{2}-\bar{q}^{2}\right)$ (9b)
$A_{3}^{\text{BH}}$ $=-\frac{2\bar{p}\bar{p}_{0}\sin\theta\sin\theta_{0}\cos\phi}{\left(\bar{E}-\bar{p}\cos\theta\right)\left(\bar{E}_{0}-\bar{p}_{0}\cos\theta_{0}\right)}\left(4\bar{E}\bar{E}_{0}-\bar{q}^{2}\right)$ (9c)
$A_{4}^{\text{BH}}$ $=\frac{2\bar{k}^{2}\left(\bar{p}^{2}\sin^{2}\theta+\bar{p}_{0}^{2}\sin^{2}\theta_{0}-2\bar{p}\bar{p}_{0}\sin\theta\sin\theta_{0}\cos\phi\right)}{\left(\bar{E}-\bar{p}\cos\theta\right)\left(\bar{E}_{0}-\bar{p}_{0}\cos\theta_{0}\right)}.$ (9d)
The last two terms may be re-organized according to Ref. [23] for further analytical tractability
$A_{3}^{\text{BH}}$ $=2\bar{k}^{2}\frac{\bar{p}^{2}\sin^{2}\theta+\bar{p}_{0}^{2}\sin^{2}\theta_{0}}{\left(\bar{E}-\bar{p}\cos\theta\right)\left(\bar{E}_{0}-\bar{p}_{0}\cos\theta_{0}\right)}$ (10)
$A_{4}^{\text{BH}}$ $=-\frac{2\bar{p}\bar{p}_{0}\sin\theta\sin\theta_{0}\cos\phi}{\left(\bar{E}-\bar{p}\cos\theta\right)\left(\bar{E}_{0}-\bar{p}_{0}\cos\theta_{0}\right)}\left(4\bar{E}\bar{E}_{0}-\bar{q}^{2}+2\bar{k}^{2}\right)$
$=-\frac{2\bar{p}\bar{p}_{0}\sin\theta\sin\theta_{0}\cos\phi}{\left(\bar{E}-\bar{p}\cos\theta\right)\left(\bar{E}_{0}-\bar{p}_{0}\cos\theta_{0}\right)}\left(2\bar{E}^{2}+2\bar{E}_{0}^{2}-\bar{q}^{2}\right),$ (11)
since $2\left(\bar{E}_{0}-\bar{E}\right)^{2}=2\bar{E}_{0}^{2}+2\bar{E}^{2}-4\bar{E}\bar{E}_{0}=2\bar{k}^{2}$ or $2\bar{k}^{2}+4\bar{E}\bar{E}_{0}=2\bar{E}_{0}^{2}+2\bar{E}^{2}$. As momenta and energies are expressed in relativistic units, the total energies of the incoming and outgoing electrons respectively are given by the relations
$\bar{E}_{0}^{2}$ $=\bar{p}_{0}^{2}+1$ (12)
$\bar{E}^{2}$ $=\bar{p}^{2}+1,$ (13)
where $\bar{E}_{0}=\bar{E}+\bar{k}$ as a result of the energy conservation.
We would like to point out that for certain combinations of energies, momenta and angles, the expressions (9a)-(9d) exhibit near exact cancellations, and careful numerical treatment is required when evaluating the expression as is. This issue has been discussed in detail by Maximon et al. [30], who proposed a refactored formulation to overcome this caveat. In the present work, this difficulty does not arise in that the integrations performed for both the unscreened and screened cases do not exhibit such canceling terms.
In the unscreened limit ($F_{Z_{s^{\prime},s}}=0$), Eq. (5) can be integrated analytically over all the possible scattering directions, recalling that $d\Omega_{p}=\sin\theta d\theta d\phi$. The integration has first been done by Sauter [45], and the result is denoted 2BN in Ref. [24]. We will refer to the DDCS without screening as the Sauter (S) formula, even though it comes from integration of the Bethe–Heiler TDCS. It reads
$\left(\frac{d^{2}\sigma_{e-i}^{Z_{s}}}{d\bar{k}d\Omega_{k}}\right)_{\text{S}}^{\text{us}}$ $=\int\left(\frac{d^{3}\sigma_{e-i}^{Z_{s}}}{d\bar{k}d\Omega_{k}d\Omega_{p}}\right)_{\text{BH}}\!d\Omega_{p}$ (14)
$=\alpha r_{e}^{2}\frac{Z_{s}^{2}}{8\pi}\frac{\bar{p}}{\bar{k}\bar{p}_{0}}\sum_{n=1}^{n=10}B_{n}^{\text{S}},$ (15)
@@
$B_{1}^{\text{S}}$ $=\frac{8\sin^{2}\theta_{0}}{\bar{p}_{0}^{2}\Delta_{0}^{4}}\left(2\bar{E}_{0}^{2}+1\right)$
$B_{2}^{\text{S}}$ $=-\frac{2\left(5\bar{E}_{0}^{2}+2\bar{E}\bar{E}_{0}+3\right)}{\bar{p}_{0}^{2}\Delta_{0}^{2}}$
$B_{3}^{\text{S}}$ $=-\frac{2\left(\bar{p}_{0}^{2}-\bar{k}^{2}\right)}{Q^{2}\Delta_{0}^{2}}$
$B_{4}^{\text{S}}$ $=\frac{4\bar{E}}{\bar{p}_{0}^{2}\Delta_{0}}$
$B_{5}^{\text{S}}$ $=\frac{L}{\bar{p}\bar{p}_{0}}\frac{4\bar{E}_{0}\sin^{2}\theta_{0}\left(3\bar{k}-\bar{p}_{0}^{2}\bar{E}\right)}{\bar{p}_{0}^{2}\Delta_{0}^{4}}$
$B_{6}^{\text{S}}$ $=\frac{L}{\bar{p}\bar{p}_{0}}\frac{4\bar{E}_{0}^{2}\left(\bar{E}_{0}^{2}+\bar{E}^{2}\right)}{\bar{p}_{0}^{2}\Delta_{0}^{2}}$
$B_{7}^{\text{S}}$ $=\frac{L}{\bar{p}\bar{p}_{0}}\frac{2-2\left(7\bar{E}_{0}^{2}-3\bar{E}\bar{E}_{0}+\bar{E}^{2}\right)}{\bar{p}_{0}^{2}\Delta_{0}^{2}}$
$B_{8}^{\text{S}}$ $=\frac{L}{\bar{p}\bar{p}_{0}}\frac{2\bar{k}\left(\bar{E}_{0}^{2}+\bar{E}\bar{E}_{0}-1\right)}{\bar{p}_{0}^{2}\Delta_{0}}$
$B_{9}^{\text{S}}$ $=-\frac{4\epsilon}{\bar{p}\Delta_{0}}$
$B_{10}^{\text{S}}$ $=\frac{\epsilon^{Q}}{\bar{p}Q}\left[\frac{4}{\Delta_{0}^{2}}-\frac{6\bar{k}}{\Delta_{0}}-\frac{2\bar{k}\left(\bar{p}_{0}^{2}-\bar{k}^{2}\right)}{Q^{2}\Delta_{0}}\right],$
@@
$L=\ln\left(\frac{\bar{E}\bar{E}_{0}-1+\bar{p}\bar{p}_{0}}{\bar{E}\bar{E}_{0}-1-\bar{p}\bar{p}_{0}}\right),$
$\Delta_{0}=\bar{E}_{0}-\bar{p}_{0}\cos\theta_{0}$, and $Q^{2}=\bar{p}_{0}^{2}+\bar{k}^{2}-2\bar{p}_{0}\bar{k}\cos\theta_{0}$. In addition,
$\epsilon=\ln\left(\frac{\bar{E}+\bar{p}}{\bar{E}-\bar{p}}\right)\quad\text{and}\quad\epsilon^{Q}=\ln\left(\frac{Q+\bar{p}}{Q-\bar{p}}\right).$ (16)
The first Born approximation, and thus the Bethe–Heitler–Sauter cross section of Eq. (15), is valid only when $\alpha Z_{s}\!\ll\!(\beta_{0},\beta)$, a very restrictive condition. Under this approximation, the Born formulas yield cross sections larger than experimental values in the soft-photon limit ($k\!\ll\!E$), but fall below them in the tip (high frequency) region, where the predicted cross section goes to zero as the photon energy reaches the incident energy.

### v2

@@ I Introduction
A recurring limitation of analytic bremsstrahlung treatments is the incorporation of screening. Classical prescriptions (Molière parameterizations [35], Thomas–Fermi–like screening lengths [53, 9], and tabulated atomic form factors [19, 46]) describe neutral-atom behavior but do not by themselves provide a compact analytic route to interpolate across charge states. Alternative analytic screening laws have also been proposed, such as the Thomas–Fermi–Kirillov form (TFK), which replaces the Yukawa exponential by a power $n=3/2$ [25], see Eq. (27), but these remain difficult to integrate consistently into bremsstrahlung cross section formalisms. Model-potential approaches, most notably that of Avdonina and Pratt [2], provide analytic screened atomic form factors (AFF) and explicitly address ionic charge states, demonstrating that screening significantly alters spectral and polarization observables. Their framework, however, relies on a specific one-parameter potential and is not easily generalized to reproduce high-quality tabulations or to interpolate smoothly between neutral and bare-ion limits. At the same time, investigations of Coulomb corrections emphasize that screening and higher-order distortion must be treated together to achieve percent-level accuracy in the application-relevant $\sim$1–100 MeV window and for moderate to high-$Z$ targets [45, 31].
To overcome these difficulties, we introduce an analytic AFF representation built from a finite sum of Yukawa screening potentials [18, 48]. Each term admits a simple analytic momentum-space component suitable for insertion into differential cross sections, while the linear combination can be fitted to high-quality atomic data (Dirac–Hartree–Fock–Slater (DHFS) fits [46, 12, 11] or Hubbell tabulations [19]) and adjusted to account for removed bound charge, thereby extending from neutral atoms to fully stripped ions. Practically, screening is implemented by inserting the fitted multi-Yukawa (MY) form factor into the Bethe–Heitler (BH) differential cross section while higher-order Coulomb distortion is retained through the Roche–Ducos–Proriol (RDP) expression for the unscreened contribution [45]. Following the Olsen–Maximon–Wergeland (OMW) additivity rule [38] (see Sec. II), the screened doubly differential cross section is obtained as the sum of the unscreened expression in a pure Coulomb field, and a screening correction in the first Born approximation, such that both neutral-atom and bare-nucleus limits are preserved. The result is a compact, rapidly evaluable expression for the doubly differential cross section (DDCS) that is inexpensive compared with full PW models.
@@ II Theoretical Framework for Electron–Ion Bremsstrahlung
$M$ $:=\int\psi^{\dagger}(\mathbf{r})(\bm{\alpha}\mathbf{\cdot e^{*}})e^{-i\bm{k\cdot r}}\psi_{0}(\mathbf{r})d^{3}\mathbf{r}.$ (2)
@@
The triply-differential cross section (TDCS) in Eq. (1) involves an integral over all space of the initial and final electron wave functions. In the vicinity of the nucleus, the nuclear (Coulomb) effects are important, while further from it, the electrons bound to the nucleus may screen the potential of the latter, so that a screening correction to the pure Coulomb field cross section ought to be considered. Olsen, Maximon and Wergeland noted that for the DDCS, i.e. after integrating out the electron’s degrees of freedom, and for energies greater than the electron rest mass, the screening correction to the Coulomb cross section is additive [38]:
$\left(\frac{d^{2}\sigma_{e-i}^{Z_{s}}}{dkd\Omega_{k}}\right)^{\text{s}}\simeq\left(\frac{d^{2}\sigma_{e-i}^{Z}}{dkd\Omega_{k}}\right)_{\text{Coul}}^{\text{us}}+\left[\left(\frac{d^{2}\sigma_{e-i}^{Z_{s}}}{dkd\Omega_{k}}\right)_{\text{Born}}^{\text{s}}-\left(\frac{d^{2}\sigma_{e-i}^{Z}}{dkd\Omega_{k}}\right)_{\text{Born}}^{\text{us}}\right].$ (3)
The latter is the so-called OMW additivity rule. The term in brackets is the screening correction. The superscripts “s” and “us” mean screened and unscreened, respectively. An illustrating figure is presented at the beginning of Section IV, while a discussion on the validity of the OMW rule can be found in [20]. We now introduce the expressions we will consider for the terms from Eq. (3).
@@ II.1 Bethe–Heitler Cross Section
$\left(\frac{d^{3}\sigma_{e-i}^{Z_{s}}}{d{k}d\Omega_{k}d\Omega_{p}}\right)_{\text{BH}}=$ $\alpha\left(\frac{r_{e}}{2\pi}\right)^{2}Z^{2}\left(1-F_{Z_{s}}\left({q}\right)\right)^{2}$ (4)
$\times\frac{{p}}{{k}p_{0}{q}^{4}}A^{\text{BH}},$ (5)
with $r_{e}$ the classical electron radius, and where $F_{Z_{s}}$ is the atomic form factor (AFF). The latter is set to zero in the unscreened case, and we retain this expression for the unscreened bremsstrahlung term, i.e. the third on the right-hand side of Eq. (3). The recoil momentum transferred to the nucleus, q{q}, is given by the relation
$\bm{q}=\bm{p}-\bm{p}_{0}-\bm{k},$ (6)
such that
${q}^{2}=$ ${p}_{0}^{2}+{p}^{2}+{k}^{2}-2{p}_{0}{k}\cos\theta_{0}+2{p}{k}\cos\theta$ (7)
$-2{p}{p}_{0}\left(\cos\theta\cos\theta_{0}+\sin\theta\sin\theta_{0}\cos\phi\right),$ (8)
where $\theta_{0}$ and $\theta$ denote the angles between the emitted photon and the incoming and scattered electron momenta, respectively, and $\phi$ is the angle between the planes $(\bm{p}_{0},\bm{k})$ and $(\bm{p},\bm{k})$. Here, and from now on, $Z_{s}:=Z-N_{s}$ represents the effective ionic charge of the species, that is, the atomic number minus the number of bound electrons. The term $A^{\text{BH}}$ in Eq. (5) is given by
$A^{\text{BH}}=\sum_{n=1}^{4}A_{n}^{\text{BH}},$ (9)
@@
$A_{1}^{\text{BH}}$ $=\frac{{p}^{2}\sin^{2}\theta}{\left({E}-{p}\cos\theta\right)^{2}}\left(4{E}_{0}^{2}-{q}^{2}\right)$ (10a)
$A_{2}^{\text{BH}}$ $=\frac{{p}_{0}^{2}\sin^{2}\theta_{0}}{\left({E}_{0}-{p}_{0}\cos\theta_{0}\right)^{2}}\left(4{E}^{2}-{q}^{2}\right)$ (10b)
$A_{3}^{\text{BH}}$ $=2{k}^{2}\frac{{p}^{2}\sin^{2}\theta+{p}_{0}^{2}\sin^{2}\theta_{0}}{\left({E}-{p}\cos\theta\right)\left({E}_{0}-{p}_{0}\cos\theta_{0}\right)}$ (10c)
$A_{4}^{\text{BH}}$ $=-\frac{2{p}{p}_{0}\sin\theta\sin\theta_{0}\cos\phi}{\left({E}-{p}\cos\theta\right)\left({E}_{0}-{p}_{0}\cos\theta_{0}\right)}\left(2{E}^{2}+2{E}_{0}^{2}-{q}^{2}\right).$ (10d)
For certain combinations of energies, momenta and angles, the expressions (10a)-(10d) exhibit near exact cancellations, and careful numerical treatment is required when evaluating the expression as is. This issue has been discussed in detail by Maximon et al. [33], who proposed a refactored formulation to overcome this caveat. In the present work, this difficulty does not arise in that the integrations performed for both the unscreened and screened cases do not exhibit such canceling terms.
In the unscreened limit ($F_{Z_{s}}=0$), Eq. (5) can be integrated analytically over all the possible scattering directions, recalling that $d\Omega_{p}=\sin\theta d\theta d\phi$. The integration has first been done by Sauter [47], and the result is denoted 2BN in Ref. [26]. We will refer to the DDCS without screening as the Sauter (S) formula, even though it comes from integration of the Bethe–Heiler TDCS. It reads
$\left(\frac{d^{2}\sigma_{e-i}^{Z}}{d{k}d\Omega_{k}}\right)_{\text{S}}^{\text{us}}$ $=\int\left(\frac{d^{3}\sigma_{e-i}^{Z}}{d{k}d\Omega_{k}d\Omega_{p}}\right)_{\text{BH}}\!d\Omega_{p}$ (11)
$=\alpha r_{e}^{2}\frac{Z^{2}}{8\pi}\frac{{p}}{{k}{p}_{0}}\sum_{n=1}^{10}B_{n}^{\text{S}},$ (12)
@@
$B_{1}^{\text{S}}$ $=\frac{8\sin^{2}\theta_{0}}{{p}_{0}^{2}\Delta_{0}^{4}}\left(2{E}_{0}^{2}+1\right)$
$B_{2}^{\text{S}}$ $=-\frac{2\left(5{E}_{0}^{2}+2{E}{E}_{0}+3\right)}{{p}_{0}^{2}\Delta_{0}^{2}}$
$B_{3}^{\text{S}}$ $=-\frac{2\left({p}_{0}^{2}-{k}^{2}\right)}{Q^{2}\Delta_{0}^{2}}$
$B_{4}^{\text{S}}$ $=\frac{4{E}}{{p}_{0}^{2}\Delta_{0}}$
$B_{5}^{\text{S}}$ $=\frac{L}{{p}{p}_{0}}\frac{4{E}_{0}\sin^{2}\theta_{0}\left(3{k}-{p}_{0}^{2}{E}\right)}{{p}_{0}^{2}\Delta_{0}^{4}}$
$B_{6}^{\text{S}}$ $=\frac{L}{{p}{p}_{0}}\frac{4{E}_{0}^{2}\left({E}_{0}^{2}+{E}^{2}\right)}{{p}_{0}^{2}\Delta_{0}^{2}}$
$B_{7}^{\text{S}}$ $=\frac{L}{{p}{p}_{0}}\frac{2-2\left(7{E}_{0}^{2}-3{E}{E}_{0}+{E}^{2}\right)}{{p}_{0}^{2}\Delta_{0}^{2}}$
$B_{8}^{\text{S}}$ $=\frac{L}{{p}{p}_{0}}\frac{2{k}\left({E}_{0}^{2}+{E}{E}_{0}-1\right)}{{p}_{0}^{2}\Delta_{0}}$
$B_{9}^{\text{S}}$ $=-\frac{4\epsilon}{{p}\Delta_{0}}$
$B_{10}^{\text{S}}$ $=\frac{\epsilon^{Q}}{{p}Q}\left[\frac{4}{\Delta_{0}^{2}}-\frac{6{k}}{\Delta_{0}}-\frac{2{k}\left({p}_{0}^{2}-{k}^{2}\right)}{Q^{2}\Delta_{0}}\right],$
@@
$L=\ln\left(\frac{{E}{E}_{0}-1+{p}{p}_{0}}{{E}{E}_{0}-1-{p}{p}_{0}}\right),$
$\Delta_{0}={E}_{0}-{p}_{0}\cos\theta_{0}$, and $Q^{2}={p}_{0}^{2}+{k}^{2}-2{p}_{0}{k}\cos\theta_{0}$. In addition,
$\epsilon=\ln\left(\frac{{E}+{p}}{{E}-{p}}\right)\quad\text{and}\quad\epsilon^{Q}=\ln\left(\frac{Q+{p}}{Q-{p}}\right).$ (13)
The first Born approximation, and thus the Bethe–Heitler–Sauter cross section of Eq. (12), is valid only when $\alpha Z\!\ll\!(\beta_{0},\beta)$, a very restrictive condition. Under this approximation, the Born formulas yield cross sections larger than experimental values in the soft-photon limit ($k\!\ll\!E$), but fall below them in the tip (high frequency) region, where the predicted cross section goes to zero as the photon energy reaches the incident energy.